OpenStudy (anonymous):

if you get 4 points for each correct answer and 1 point for each wrong answer and earned 50 points by answering 25 questions what would the equation be?

OpenStudy (anonymous):

do you mean one point taken off for each wrong answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x - y = 50, x + y = 25, where x is the number of correct answers and y is the number of wrong answers. Then solve as you would simultaneous equasions: 4x - y = 50 x + y = 25 5x = 75 x = 15 thefore if x + y = 25, y =10

OpenStudy (anonymous):

thanks.. i was trying to put in one equation.